# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v1

@@ 4 Detection in correlated SBMs
The crucial input in our proof of Lemma 4.4 is the following estimation on $\Xi(H)$.
Lemma 4.5.
@@
${}|\Xi(C)|\leq k(1-\delta)^{\frac{|E(C)|}{2}}n^{-\frac{|E(C)|}{2}}\,.$ (4.23)
@@
Now we prove Lemma 4.4 based on Lemma 4.5.
@@
$\sum_{\sigma\in[k]^{n}}\sum_{H\Subset\mathcal{K}_{n},|E(H)|\leq D}u_{\sigma,H}^{2}=\sum_{\begin{subarray}{c}H\Subset\mathcal{K}_{n},\mathcal{L}(H)=\emptyset\\ |E(H)|\leq D\end{subarray}}\Xi(H)^{2}$ (4.24)
is bounded by $O_{\delta}(1)$. Using Lemma 4.5, we see that (4.24) is bounded by
$\sum_{\begin{subarray}{c}m_{3},\ldots,m_{D}\geq 0,t\geq 0\\ v\geq 3m_{3}+\ldots+Dm_{D}\end{subarray}}2^{10k+m_{3}+\ldots+m_{D}}k^{10t}n^{-(v+t)}(1-\delta/2)^{v+t}\cdot\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)\,,$ (4.25)
@@
${}\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)=\#\Big{\{}S\Subset\mathcal{K}_{n}:\mathcal{L}(S)=\emptyset,\mathfrak{C}_{k}(S)=m_{k}\mbox{ for }k\leq D,\tau(S)=t\Big{\}}\,.$ (4.26)
@@
${}\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)\leq n^{v}D^{10t}\prod_{3\leq i\leq D}\frac{1}{m_{i}!}\,.$ (4.27)
Plugging (4.27) into (4.25), we get that
@@
$\leq O_{\delta}(1)\cdot\sum_{m_{3},\ldots,m_{D}\geq 0}2^{m_{3}+\ldots+m_{D}}\prod_{3\leq i\leq D}\frac{1}{i^{m_{i}}m_{i}!}=O_{\delta}(1)\,.$
@@
Acknowledgment. The author thanks Hang Du for pointing him to the reference [52]; the author thanks Jingqiu Ding for stimulating discussions on the strengthening of low-degree conjecture. This work is partially supported by National Key R&\&&D program of China (No. 2023YFA1010103) and NSFC Key Program (Project No. 12231002).
Appendix A Preliminary results in graphs and probability
In this section we include some preliminary results we will use in the main part of the paper. We first state some results on the structural properties of graphs established in [15, 9].
Lemma A.1 ([15], Lemma A.1).
  1. (i)
  2. (ii)
    $\Phi(S\cup T)\Phi(S\Cap T)\leq\Phi(S)\Phi(T)$.
  3. (iii)
  4. (iv)
  5. (v)
Lemma A.2.
Let $S,T\subset\mathcal{K}_{n}$. Recall that $S\Cap T\Subset\mathcal{K}_{n}$ is defined as edge-induced subgraphs of $\mathcal{K}_{n}$. We have the following properties:
  1. (i)
    $\tau(S\cup T)+\tau(S\Cap T)\geq\tau(S)+\tau(T)$ and $\Upsilon(S\cup T)\Upsilon(S\Cap T)\leq\Phi(S)\Phi(T)$.
  2. (ii)
    $|\mathtt{C}_{j}(S\cup T)|+|\mathtt{C}_{j}(S\cap T)|\geq|\mathtt{C}_{j}(S)|+|\mathtt{C}_{j}(T)|$.
  3. (iii)
    If $S\subset T$, $S$ is self-bad and $V(S)=V(T)$, then $T$ is self-bad.
  4. (iv)
    If $S$ and $T$ are both self-bad, then $S\cup T$ is self-bad.
Lemma A.3 ([9], Lemma A.3).
@@ Appendix A Preliminary results in graphs and probability
  3. (3)
  4. (4)
    $\mathtt{t}=|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H)$.
Lemma A.4 ([9], Corollary A.4).
  1. (1)
    ${C}_{\mathtt{1}},\ldots,{C}_{\mathtt{m}}$ are independent cycles in $S$.
  2. (2)
  3. (3)
    $\mathtt{t}\leq 5(|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H))$.
Lemma A.5 ([9], Lemma A.9).
Lemma A.6 ([9], Lemma A.10).
For $S\subset\mathcal{K}_{n}$ with $|E(S)|\leq D$, we have
@@
Lemma A.7.
Proof.
It remains to prove that
which completes the induction procedure. ∎
Lemma A.8.
For $H\subset S$ with $\mathcal{L}(S)\not\subset V(H)$, we have
Proof.
=\ = $\mathbb{E}_{\sigma_{[n]\setminus\mathtt{V}}\sim\nu_{\mathtt{V}}}\Big{[}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\Big{]}$
=\ = $\mathbb{E}_{\sigma_{[n]\setminus(\mathtt{V}\cup\{v\})}\sim\nu_{\mathtt{V}}}\mathbb{E}_{\sigma_{\{v\}}\sim\nu_{\{v\}}}\Big{[}\omega(\sigma_{v},\sigma_{w})\prod_{(i,j)\in E(S)\setminus(E(H)\cup\{v,w\})}\omega(\sigma_{i},\sigma_{j})\Big{]}=0\,.\qed$

### v2

@@ Appendix B Supplementary proofs
In this subsection we present the proof of Lemma 4.4. The crucial input in our proof of Lemma 4.4 is the following estimation on $\Xi(H)$.
Lemma B.5.
@@
${}|\Xi(C)|\leq k(1-\delta)^{\frac{|E(C)|}{2}}n^{-\frac{|E(C)|}{2}}\,.$ (B.43)
@@
Now we prove Lemma 4.4 based on Lemma B.5.
@@
$\sum_{\sigma\in[k]^{n}}\sum_{H\Subset\mathcal{K}_{n},|E(H)|\leq D}u_{\sigma,H}^{2}=\sum_{\begin{subarray}{c}H\Subset\mathcal{K}_{n},\mathcal{L}(H)=\emptyset\\ |E(H)|\leq D\end{subarray}}\Xi(H)^{2}$ (B.44)
is bounded by $O_{\delta}(1)$. Using Lemma B.5, we see that (B.44) is bounded by
$\sum_{\begin{subarray}{c}m_{3},\ldots,m_{D}\geq 0,t\geq 0\\ v\geq 3m_{3}+\ldots+Dm_{D}\end{subarray}}2^{10k+m_{3}+\ldots+m_{D}}k^{10t}n^{-(v+t)}(1-\delta/2)^{v+t}\cdot\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)\,,$ (B.45)
@@
${}\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)=\#\Big{\{}S\Subset\mathcal{K}_{n}:\mathcal{L}(S)=\emptyset,\mathfrak{C}_{k}(S)=m_{k}\mbox{ for }k\leq D,\tau(S)=t\Big{\}}\,.$ (B.46)
@@
${}\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)\leq n^{v}D^{10t}\prod_{3\leq i\leq D}\frac{1}{m_{i}!}\,.$ (B.47)
Plugging (B.47) into (B.45), we get that
@@
$\leq O_{\delta,k}(1)\cdot\sum_{m_{3},\ldots,m_{D}\geq 0}2^{m_{3}+\ldots+m_{D}}\prod_{3\leq i\leq D}\frac{1}{i^{m_{i}}m_{i}!}=O_{\delta,k}(1)\,.$
@@
Appendix C Proof of Proposition 3.3 assuming the low-degree conjecture
Shortly after the submission of an earlier version of this draft, we realized that actually we can prove Proposition 3.3 (and thus proving Theorem 3.2) assuming only the original low-degree conjecture. However, such a proof seems to be more problem dependent so we postpone it here. To be more precise, in this section we will prove Proposition 3.3 assuming the following (truncated) low-degree conjecture.
Conjecture C.1 (Truncated low-degree conjecture).
@@
Our proof similarly starts by Lemma 3.5. The key observation is that the following weaker version of Lemma 3.6 suffices to deduce the impossibility of partial matching.
Lemma C.2.
Assuming Conjecture C.1, we have for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ such that each $g_{i,j}(A,B)$ can be computed in time $n^{\widetilde{O}(D)}$ and
$\sum_{1\leq j\leq n}g_{i,j}(A,B)=o(n)\mbox{ for all }1\leq i\leq n\mbox{ and for all graphs }A,B\mbox{ on }[n]\,,$ (C.1)
it holds that
${}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{i,j}\big{)}^{2}\Big{]}\geq 1-o(1)\,.$ (C.2)
Proof.
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$ such that $\sum_{1\leq j\leq n}g_{j}(A,B)=o(n)$ for all graphs $A,B$ and (3.6) holds. Again we may assume that $0\leq g_{i}\leq 1$. Recall (3.7), (3.8) and (3.9). Using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture C.1) we have (3.10) still holds. To this end, define $g=g_{1}+\ldots+g_{n}$. Using $0\leq g\leq n$, we see that (3.10) yields that $\mathbb{Q}(g\geq\Omega(n))\geq\Omega(n)$, which violates our assumption (C.1). This implies that (3.6) is impossible and thus completes our proof. ∎
Now we can finish the proof of Item (1) in Proposition 3.3 assuming Conjecture C.1.
Proof of Item (1) in Proposition 3.3 assuming Conjecture C.1.
Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$ satisfying Items (1) and (2) in Definition 3.1 with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. We first claim that we may assume that
$f_{i,j}(A,B)\in\{0,1\}\mbox{ and }\sum_{1\leq j\leq n}f_{i,j}(A,B)\in\{0,1\}\mbox{ for all graphs }A,B\mbox{ on }[n]\,.$ (C.3)
In fact, denote
$\mathcal{A}=\Big{(}\cap_{1\leq i,j\leq n}\Big{\{}f_{i,j}(A,B)\in\{0,1\}\Big{\}}\Big{)}\bigcap\Big{(}\cap_{1\leq i\leq n}\Big{\{}\sum_{1\leq j\leq n}f_{i,j}(A,B)=1\Big{\}}\Big{)}\,,$
we may consider the truncated statistics
$f_{i,j}^{\prime}(A,B)=f_{i,j}(A,B)\cdot\mathbf{1}_{\mathcal{A}}\,.$
It is straightforward to check that $\{f_{i,j}^{\prime}(A,B):1\leq i,j\leq n\}$ satisfies (C.3) and we still have
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{i,j}^{\prime}(A,B)\big{]}\geq 1-c\,.$
Thus, we may assume that (C.3) holds without losing of generality. To this end, using (C.3), we see that $f_{j}f_{k}=0$ for all $k\neq j$, and thus
$1=\mathbb{E}_{\mathbb{P}}\big{[}(f_{1}+\ldots+f_{n})^{2}\big{]}=\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}}\big{[}f_{j}^{2}\big{]}\,.$ (C.4)
In addition, we have
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{\pi_{*}(i)}\big{]}=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{j}\mid\pi_{*}(i)=j\big{]}\geq c\,.$ (C.5)
Thus, for all $\lambda\in[0,1]$ we have
@@
contradicting to Lemma 3.6. This leads to the desired result. ∎